### v1

# DKDS: A Benchmark Dataset of Degraded Kuzushiji Documents with Seals for Detection and Binarization

Abstract
Kuzushiji, a pre-modern Japanese cursive script, can currently be read and understood by only a few thousand trained experts in Japan. With the rapid development of deep learning, researchers have begun applying Optical Character Recognition (OCR) techniques to transcribe Kuzushiji into modern Japanese. Although existing OCR methods perform well on clean pre-modern Japanese documents written in Kuzushiji, they often fail to consider various types of noise, such as document degradation and seals, which significantly affect recognition accuracy. To the best of our knowledge, no existing dataset specifically addresses these challenges. To address this gap, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset as a new benchmark for related tasks. We describe the dataset construction process, which required the assistance of a trained Kuzushiji expert, and define two benchmark tracks: (1) text and seal detection and (2) document binarization. For the text and seal detection track, we provide baseline results using multiple versions of the You Only Look Once (YOLO) models for detecting Kuzushiji characters and seals. For the document binarization track, we present baseline results from traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods. The DKDS dataset and the implementation code for baseline methods are available at https://ruiyangju.github.io/DKDS.
keywords:
Kuzushiji, Seals, Pre-modern Japanese Documents, Benchmark Dataset, Object Detection, Document Binarization
1 Introduction
Pre-modern Japanese documents were mainly written in Kuzushiji111くずし字, a form of Japanese cursive script (草書体)., a pre-modern Japanese cursive script widely used in official records, personal letters, and literary works. Kuzushiji is characterized by complex, flowing strokes and highly variable forms, which differ markedly from modern Japanese characters [ueki2021survey]. Educational reforms in the early 20th century introduced standardized textbooks and led to reduced instruction in Kuzushiji. Following the adoption of modern kana orthography in 1946, most contemporary native Japanese speakers are unable to read Kuzushiji texts without specialized training [hashimoto2017Kuzushiji].
Figure 1: Comparison of Optical Character Recognition (OCR) results on Kuzushiji characters overlapping with seals between the original and binarized images. OCR was conducted using the “miwo” app [clanuwat2021miwo]. From left to right, the observed OCR errors include recognition of extra character, recognition of incorrect character, and misclassification of seal inscriptions as text.
With the remarkable advances in deep learning for computer vision and natural language processing, researchers have increasingly applied neural network models to Kuzushiji character recognition [le2019human, le2020automated]. While trained models typically achieve satisfactory performance in recognizing Kuzushiji characters seen during training, their performance often drops significantly when tested on unseen (zero-shot) Kuzushiji characters. For instance, the framework proposed by Ishikawa et al. [ishikawa2024japanese] achieved an accuracy of only approximately 48% in such zero-shot scenarios. This result presents the critical need to construct high-quality Kuzushiji datasets, including special cases, to improve model generalization and capabilities in complex recognition tasks.
The primary Kuzushiji datasets include Kuzushiji-MNIST, Kuzushiji-49, and Kuzushiji-Kanji [^2] Kanji refers to Chinese-origin characters used in Japanese. [clanuwat2018deep]. The Kaggle competitions organized using these datasets has further advanced both research and practical applications [kitamoto2019progress, kitamoto2020kaggle]. Together, these datasets and competitions cover a wide range of Kuzushiji characters and provide a standardized evaluation platform for researchers. With these resources, researchers can perform not only Kuzushiji character classification but also related tasks such as Kuzushiji detection and generation [soan2023], thereby advancing the development of Kuzushiji character recognition techniques and enhancing model performance.
Several methods have been proposed for the Kuzushiji character recognition task. Fujita et al. [fujita2024learning] employed Transformer [vaswani2017attention] models to analyze and recognize Kuzushiji characters, leveraging attention mechanisms to capture relationships between adjacent characters and accurately separate character regions and boundaries. Ngo et al. [ngo2021recurrent] introduced RNN-Transducer [graves2012sequence] models for handwritten Kuzushiji character recognition, effectively exploiting both visual and text information from the input images and achieving satisfactory performance.
Figure 2: DKDS dataset is the first collection of degraded pre-modern Japanese document images specifically designed to address the challenge of Kuzushiji characters overlapping with seals. Based on the dataset, we define two benchmark tracks: (1) Text and Seal Detection, and (2) Document Binarization.
In addition, KuroNet [clanuwat2019kuronet, lamb2020kuronet] is an end-to-end model based on a residual U-Net [ronneberger2015u] architecture, designed to capture long-range context, process extensive vocabularies, and accommodate non-standard character layouts, thereby enabling the recognition of pre-modern Japanese documents. On the Kuzushiji dataset provided by the Kaggle competitions [kitamoto2019progress, kitamoto2020kaggle], KuroNet achieved an accuracy exceeding 90%. To facilitate the deployment of OCR models on mobile devices, Clanuwat et al. [clanuwat2021miwo] developed the “miwo” application, which allows users to upload or photograph Kuzushiji documents and automatically transcribe the characters into modern Japanese.
The aforementioned OCR methods and applications perform well on clean pre-modern Japanese documents written in Kuzushiji; however, they do not sufficiently consider the effects of noise, such as document degradation and seals, on recognition accuracy. As shown in Fig. 1, such noise can cause errors that significantly reduce recognition accuracy. To address this issue, document binarization, which preserves the foreground (black) while removing background and noise (white), has been widely adopted as a preprocessing step, as it can effectively improve OCR performance.
Among the various types of noise, the overlap between Kuzushiji characters and seals poses a major challenge. Serving as symbols of ownership and collector identity, seals were widely used in ancient and pre-modern Asian document collections [li2021character]. They are typically inscribed with stylized ancient scripts and may contain not only the collector’s name but also phrases expressing personal wishes, interests, or social status. In ancient and pre-modern documents, seals usually appear as red marks that frequently overlap with characters, which can blur the characters and lead to OCR errors.
To address this challenge, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset, developed in collaboration with a trained Kuzushiji expert. This novel benchmark dataset comprises degraded pre-modern Japanese documents written in Kuzushiji, with various types of seals randomly added to simulate common types of interference. As shown in Fig. 2, we define two task tracks for the DKDS dataset: (1) text and seal detection, and (2) document binarization. The former track aims to accurately detect text (i.e., Kuzushiji characters) and seal regions within the documents, while the latter focuses on preserving the primary textual content while removing noise, including stains and seals, thereby providing clean input for subsequent OCR tasks.
Overall, the contributions of this work are summarized as follows:
  • (a)
    We introduce a novel benchmark dataset, named Degraded Kuzushiji Documents with Seals (DKDS), developed in collaboration with a trained Kuzushiji expert. To the best of our knowledge, this is the first publicly available dataset specifically addressing the challenge of Kuzushiji characters overlapped with seals.
  • (b)
    We provide a detailed description of the dataset construction process, including the generation of binarization ground-truth, its verification and manual correction, and the random addition of seals to the raw Kuzushiji documents.
  • (c)
    We define the text and seal detection task, which serves as a preliminary step for recognizing Kuzushiji characters and transcribing them into modern Japanese. Baseline results for this task are also presented using multiple versions of the YOLO models.
  • (d)
    We define the document binarization task specifically for pre-modern Japanese document images. Compared to existing binarization datasets, the DKDS dataset presents additional challenges due to the overlap of Kuzushiji characters with seals. Baseline results for this task are provided, including traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods.
The rest of this paper is organized as follows: Section 2 reviews related datasets and previous work relevant to the two task tracks defined here, highlighting their limitations. Section 3 describes the dataset construction process, including raw data collection, generation of binarization ground-truth, and the procedure for adding seals. Section 4 defines the two task tracks and their corresponding evaluation metrics, and introduces the baseline methods used in this work. Section 5 presents the experimental setup and implementation details, followed by a comparative analysis of quantitative results for the two tracks. Finally, Section 6 summarizes this work and discusses the direction of future research.
2 Related Works
2.1 Related Dataset
The datasets relevant to this work can be categorized into three domains: (1) Kuzushiji document datasets, (2) degraded pre-modern document binarization datasets, and (3) seal datasets.
Existing research on pre-modern Japanese documents written in Kuzushiji [clanuwat2018deep, le2019human] lacks a dataset that specifically addresses the challenge of overlaps between characters and seals, which can significantly reduce recognition accuracy.
Most publicly available datasets for degraded ancient document binarization focus on documents written in Latin-based alphabetic systems. For instance, the documents in the Document Image Binarization Contest (DIBCO) series datasets [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2012icfhr, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019] are primarily written in English, while the Bickley Diary (BD) dataset [deng2010binarizationshop], Synchromedia Multispectral Ancient Document Images (SMADI) [hedjam2013historical], and Persian Heritage Image Binarization Dataset (PHIBD) [ayatollahi2013persian] all contain documents written in European languages. In contrast, no publicly available document binarization datasets exist for pre-modern Japanese documents, particularly those written in Kuzushiji.
Furthermore, no existing degraded document datasets include seals as the primary source of interference. Therefore, we aim to randomly add seals to the collected pre-modern Japanese documents written in Kuzushiji. However, existing publicly available seal datasets are all unsuitable, as the seals they contain are either not ancient Asian seals, suffer from poor image quality, or have complex backgrounds. Specifically, the Chinese-Seal Dataset (CSD) [xu2025chinese] focuses on modern Chinese seals; DDI-100 [zharikov2020ddi] contains Western (English-language) seal samples; and SPODS [nandedkar2016spods] comprises seal images from contemporary office settings. Although MiikeMineStamps [buitrago2021miikeminestamps] includes some Japanese seal images, they are generally low-resolution and feature non-uniform backgrounds, making background removal difficult.
2.2 Text and Seal Detection
The diverse shapes of ancient seals, their curved inscriptions, and the frequent occurrence of text overlap pose significant challenges for seal processing and text analysis and recognition. As a critical preliminary step for Kuzushiji character recognition, text detection is particularly important [li2023yolo]. Similarly, seal detection is important for subsequent seal processing.
However, the scarcity of ancient seal samples and the prevalence of low-quality or degraded images severely limit available training data and annotations for related tasks [li2021character]. For instance, Micenkov et al. [micenkov2011stamp] proposed an automatic seal segmentation system for insurance documents that achieved 84% accuracy. Yu et al. [yu2023icdar] developed a modern seal dataset for official and financial contexts, introducing two tasks: seal title text detection and end-to-end seal title recognition. To the best of our knowledge, no dataset currently exists for text and seal detection in degraded pre-modern Japanese Kuzushiji documents.
The YOLO series of models [jocher2023yolov8, wang2024yolov9, wang2024yolov10, jocher2024yolo11] achieve an excellent balance between detection accuracy and model size, demonstrating outstanding performance across a wide range of object detection tasks [ju2023fracture, chien2025yolov8, chien2024yolov9, ju2024yolov8, ju2024global]. Researchers have explored the application of the YOLO series of models in Kuzushiji character detection [li2023yolo] and seal detection [bento2025performance]. For instance, Bento et al. [bento2025performance] merged and adapted the StaVer [micenkov2011stamp] and DDI-100 [zharikov2020ddi] datasets to evaluate the performance of YOLOv8, YOLOv9, YOLOv10, and YOLO11 on their newly constructed dataset for seal detection. Similarly, Sun et al. [sun2024ra] enhanced the YOLOv8 architecture and proposed RA-YOLOv8, achieving higher detection accuracy. Nevertheless, due to the lack of seal data based on Kuzushiji documents, these improved YOLO models cannot be directly applied to this specific scenario.
Figure 3: Examples of raw Kuzushiji document data from the book Genji Monogatari (The Tale of Genji[genjimonogatari].
2.3 Document Binarization
Document binarization is an important preprocessing step for OCR and layout analysis, effectively removing various types of noise from document images, such as paper yellowing, text fading, ink bleeding, stains, and seals. The Document Image Binarization Contest (DIBCO), one of the most popular competitions in this field, has provided benchmark datasets from 2009 to 2019 [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2012icfhr, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019]. Researchers have proposed a range of neural network models evaluated on these datasets, including those based on GANs [suh2022two, ju2023ccdwt, ju2024mfe], Transformers [souibgui2022docentr], and diffusion models [saifullah2023coldbin, yang2023docdiff], which have significantly advanced the field. However, most existing datasets contain ancient and pre-modern document images based on Latin alphabets, while few include Kuzushiji documents, particularly those containing overlapping seals and text.
Figure 4: Examples of our collected imperial seals from the Qing dynasty, with backgrounds removed.
Figure 5: The overall workflow of the proposed DKDS dataset construction includes raw data collection, detection annotations, initial binarization ground-truth generation, verification, and manual correction. The initial binarization ground-truth was generated using the model proposed by Ju et al. [ju2024three], while the verification was conducted by a trained Kuzushiji expert.
3 Dataset Construction
3.1 Raw Data Collection
The raw Kuzushiji document data were obtained from the Center for Open Data in the Humanities (CODH) [genjimonogatari]. Specifically, we selected the dataset corresponding to book ID “200003803”, titled Genji Monogatari (The Tale of Genji) [^3] https://codh.rois.ac.jp/char-shape/book/200003803/. Examples from this book are shown in Fig 3. This dataset was chosen for two main reasons: (1) its historical significance and widespread popularity; and (2) the availability of detailed OCR annotations provided by CODH, including 237 unique character types and a total of 11,132 characters. These data can be used for training and evaluating OCR models in future research.
Furthermore, to date, there is no publicly available high-resolution dataset of ancient Japanese seals. To address this gap, we collected 219 seals from three emperors of the Qing dynasty (i.e., Kangxi (1661–1722), Yongzheng (1722–1735), and Qianlong (1735–1796)) sourced from the developer of an ancient seal app [^4] https://www.daizigege.com/seal/. The primary reason for selecting ancient Chinese seals is that both ancient Japanese and Chinese seals are mainly inscribed with kanji characters.
Considering factors such as image resolution and seal patterns (e.g., whether the inscriptions are written in kanji), we selected a final set of 128 seal images from the collection. As shown in Fig. 4, these seals are high-resolution, and their backgrounds can be easily removed.
Table 1: Statistics of the proposed DKDS dataset.
Category Total Training Testing
Kuzushiji Documents 50 40 10
Seals 128 $100\times 2$ 28
Seals per Document - 4 – 6 2 – 3
Figure 6: Examples of verification and manual correction in the dataset construction workflow: “Raw” refers to the input images, “Initial” denotes the initial binarization produced by Ju et al. [ju2024three], and “Corrected” represents the final ground-truth after verification and manual correction. The trained Kuzushiji expert performed verification to determine which elements should be preserved (e.g., marginal notes, voiced sound symbols, periods) and which should be removed (e.g., stains).
Figure 7: Challenging examples for the text and seal detection task: The first row presents seals suffering from ink fading, while the second row shows seals overlapping with Kuzushiji characters or other seals.
3.2 Construction Workflow
We constructed our dataset by producing color document images and their corresponding binarized ground-truth images from the raw data, and the overall workflow is shown in Fig. 5. To create the final document images, we combined 128 background removed seal images with 50 Kuzushiji document images. Specifically, 40 Kuzushiji documents were selected for training, and the remaining 10 for testing. As detailed in Table 1, the 128 seal images were divided into 100 for the 40 training Kuzushiji documents, and 28 for the 10 test Kuzushiji documents. To increase the cases of overlap between seals and Kuzushiji characters in the training set, we duplicated the 100 training seal images to create 200 seals, which were then randomly added onto the 40 training Kuzushiji documents. As a result, each Kuzushiji document image in the training set contains approximately 4–6 seals, while those in the test set contain about 2–3 seals.
To generate the binarization ground-truth, we employed a combined approach of neural network generation followed by manual verification and correction. As shown in Fig. 5, we first applied the model proposed by Ju et al. [ju2024three] to binarize the raw Kuzushiji document images (without added seals), producing the initial binarization ground-truth. Subsequently, a trained Kuzushiji expert manually verified these initial results. Even in the absence of seal interference, the initial binarization often contained errors. As detailed in Fig. 6, the expert identified issues such as missing marginal notes, voiced sound symbols, periods, and unremoved stains. These errors were then manually corrected to create the final binarization ground-truth. Notably, this verification step was critical and required a trained Kuzushiji expert to ensure the validity of the process.
3.3 Detection Annotations
The detection annotations include bounding boxes for both Kuzushiji characters and seals. The bounding box information for Kuzushiji characters was obtained from the OCR annotations provided by CODH [genjimonogatari], while the bounding boxes for the seals were recorded during the process of randomly adding them to the Kuzushiji document images. Notably, the OCR annotations from CODH [genjimonogatari] include only the main Kuzushiji text, excluding marginal notes. Considering practical applications, we followed the same strategy.
All annotations are provided in the YOLO format, enabling researchers to directly download and use the dataset. In addition, detailed information for both Kuzushiji characters and seals, including their coordinates and sizes, is available on our GitHub repository, allowing researchers to convert the data into other object detection formats for model training and evaluation.
4 Task Definition
4.1 Evaluation Tracks
We define two task tracks in the proposed dataset: (1) text and seal detection, and (2) document binarization. As shown in Fig. 7, seals may overlap with Kuzushiji characters or other seals, and suffer from ink fading, all of which can significantly reduce detection accuracy, making this task highly challenging. In the document binarization task, the goal is to remove seals while preserving, or even restoring, the underlying Kuzushiji characters. This task becomes particularly difficult when the characters and seals overlap, as shown in Fig. 8.
Furthermore, both tasks hold important research significance. Text and seal detection serves as a critical preliminary step for downstream applications, including Kuzushiji character recognition and seal processing, while document binarization aims to improve the accuracy of subsequent OCR systems. By addressing these challenges, the proposed dataset encourages the development of robust models capable of processing real-world scenarios in Kuzushiji document analysis.
Figure 8: Challenging examples for the document binarization task: The left column shows the input Kuzushiji document images, while the right column presents the corresponding binarized images.
4.2 Evaluation Metrics
For the text and seal detection task, we adopted standard object detection evaluation metrics [lin2014microsoft], including model parameters (Params), floating-point operations (FLOPs), average precision at a 50% intersection over union (IoU) threshold ($\text{AP}_{50}$), and average precision across IoU thresholds from 50% to 95% ($\text{AP}_{50:95}$). Model parameters reflect the model’s size and complexity, FLOPs quantify computational cost, and these average precision (AP) metrics provide a comprehensive measure of detection accuracy.
For the document binarization task, we employed classical evaluation metrics [gatos2009icdar] for quantitative comparison, including the F-measure (FM), pseudo F-measure (p-FM), peak signal-to-noise ratio (PSNR), and distance reciprocal distortion (DRD). To provide a more comprehensive assessment, we also adopted the Average-Score Metric (ASM) proposed by Ju et al. [ju2025efficient], which is calculated as follows:
$ASM=\dfrac{FM+p\text{-}FM+PSNR+(100-DRD)}{4}.$ (1)
A higher ASM value indicates better overall performance, providing a single unified metric for model ranking.
Figure 9: The architecture of the GAN-based method used.
4.3 Baselines
For the text and seal detection task, we employed the YOLO series of object detection models, including YOLOv8 [jocher2023yolov8], YOLOv9 [wang2024yolov9], YOLOv10 [wang2024yolov10], and YOLO11 [jocher2024yolo11]. As one-stage detectors, these models reduce computational complexity and model parameters while maintaining high detection accuracy, achieving a good balance between inference speed and performance. In addition, they support end-to-end training, provide fast inference, and present strong generalization capabilities, making them widely adopted in object detection tasks.
For the document binarization task, we first applied several traditional binarization algorithms to the Kuzushiji documents, including Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive], which can be directly employed without any training.
Considering that seals in the pre-modern Japanese documents are typically red, while Kuzushiji characters are black and the paper background is pale yellow, we introduced a preprocessing step to remove seal regions based on K-Means clustering [mcqueen1967some] before applying the binarization algorithms. This preprocessing effectively reduces the interference caused by red seals and enhances the quality of subsequent binarization results. Specifically, we designed and implemented three combined approaches: K-Means + Otsu, K-Means + Niblack, and K-Means + Sauvola.
In addition, to provide a deep learning-based baseline, we employed a GAN [goodfellow2020generative]-based model to generate binarized Kuzushiji document images. The network architecture for this model is shown in Fig. 9. Since directly inputting the entire image (approximately $2100\times 3200$ pixels) would require excessive GPU memory, we divided each document image into multiple $512\times 512$ patches; the details of this division are described in Section 5.1.2. These patches were then fed into a U-Net [ronneberger2015u] architecture with an EfficientNet-B5 [tan2019rethinking] backbone, which served as the generator. A simplified PatchGAN [isola2017image, zhu2017unpaired] discriminator was adopted to distinguish between real and fake (generated) images. Furthermore, following Suh et al. [suh2022two] and Ju et al. [ju2024three], we employed a loss function that combines the Wasserstein GAN with Gradient Penalty (WGAN-GP) loss [gulrajani2017improved] and an additional Binary Cross-Entropy (BCE) loss.
Table 2: Quantitative comparison of various YOLO models for text and seal detection on the DKDS dataset.
Model Params FLOPs $\text{AP}_{50}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50}^{\text{Seal}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Seal}}$ $\uparrow$
YOLOv8m [jocher2023yolov8] 25.86M 79.1G 96.4% 71.2% 99.1% 86.2%
YOLOv9m [wang2024yolov9] 20.16M 77.5G 96.3% 71.7% 97.2% 81.4%
YOLOv10m [wang2024yolov10] 16.49M 64.0G 96.2% 71.4% 99.1% 85.7%
YOLO11m [jocher2024yolo11] 20.05M 68.2G 97.8% 74.1% 98.5% 85.7%
Figure 10: Visual comparison of text and seal detection results predicted by different models with the corresponding ground-truth annotations. The input images are “200003803_00028_1” (bottom) and “200003803_00028_2” (top).
Table 3: Quantitative comparison of various methods for document binarization on the DKDS dataset.
[…]
Niblack [niblack1985introduction] 39.13 41.14 8.44dB 79.70 27.25
Otsu [otsu1979threshold] 63.01 63.31 11.76dB 37.69 50.10
Sauvola [sauvola2000adaptive] 87.87 90.99 18.34dB 7.01 72.55
K-means [mcqueen1967some] + Niblack [niblack1985introduction] 39.99 42.03 8.61dB 76.67 28.49
K-means [mcqueen1967some] + Otsu [otsu1979threshold] 84.76 86.28 17.14dB 9.90 69.57
K-means [mcqueen1967some] + Sauvola [sauvola2000adaptive] 88.59 91.48 18.65dB 6.37 73.09
GAN [goodfellow2020generative] 98.11 98.14 26.53dB 0.82 80.49
Figure 11: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00028_2” from the DKDS test set. The input image, ground-truth, and method outputs are shown for comparison.
Figure 12: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00029_1” from the DKDS test set. The input image, ground-truth, and method outputs are shown for comparison.
5 Experiments
5.1 Implementation Details
To ensure a fair performance comparison on the DKDS dataset, all models were trained and evaluated on the same hardware (i.e., NVIDIA RTX 3090 GPU). All experiments were conducted using the Ubuntu operating system, Python 3.9, and the PyTorch framework.
5.1.1 Text and Seal Detection
For the text and seal detection task, the YOLO models were trained for 100 epochs with a batch size of 16 and an input resolution of $640\times 640$ pixels. Model training was performed using the SGD optimizer [robbins1951stochastic] with an initial learning rate of 0.01. All YOLO models were initialized with weights pre-trained on the COCO dataset [lin2014microsoft].
5.1.2 Document Binarization
For the document binarization task, traditional binarization algorithms (i.e., Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive]) first convert the input image into a grayscale image as a preprocessing step. The Niblack algorithm uses a $25\times 25$ sliding window with $k=0.8$, while the Sauvola algorithm employs the same window size. For the K-Means clustering method [mcqueen1967some], the number of clusters is set to $K=3$. After clustering, median filtering with a kernel size of 5 is applied to smooth the results and remove isolated noise points.
For the GAN-based method [goodfellow2020generative], the training data were preprocessed by dividing each original image into $512\times 512$ overlapping patches with a 30% overlap ratio. To further expand the training set, we applied data augmentation, including scaling (factors 0.75, 1.0, 1.25, and 1.5) and rotations (0°and 270°). After preprocessing and augmentation, the 40 original training images were expanded into a total of 71,458 patches of size $512\times 512$ pixels. Both the generator and discriminator were optimized using the Adam [kingma2014adam] optimizer with a learning rate of $2\times 10^{-4}$ and $\beta$ coefficients of $(0.5,0.999)$. The model was trained for 10 epochs with a batch size of 16. At inference, the predicted outputs were binarized using a 0.5 threshold.
5.2 Track 1: Text and Seal Detection
5.2.1 Quantitative Comparison
We trained and evaluated various YOLO series models on our proposed DKDS dataset, using the medium (M) size variant for all models. Quantitative comparison results are summarized in Table 2.
For the text (i.e., Kuzushiji character) detection, all models achieved over 96% in $\text{AP}_{50}$, indicating that they can accurately detect most Kuzushiji characters. Under stricter IoU thresholds (i.e., $\text{AP}_{50:95}$), YOLO11m achieved the highest score of 74.1%, outperforming the other models by about 2–3%. This demonstrates the superior accuracy of YOLO11m in localizing Kuzushiji character bounding boxes.
For the seal detection, both YOLOv8m and YOLOv10m obtained the highest $\text{AP}_{50}$ of 99.1%, indicating near-perfect detection performance under the loose IoU criterion. In addition, YOLOv8m reached the highest $\text{AP}_{50:95}$ of 86.2%, demonstrating strong localization accuracy under stricter IoU thresholds.
In terms of model size, YOLOv10m presents the lowest Params of 16.49M and FLOPs of 64.0G, making it the most lightweight model.
Meanwhile, YOLO11m achieves the best balance between accuracy and efficiency, with 20.05M parameters, 68.2G FLOPs, $\text{AP}_{50}^{Kuzushiji}$ of 97.8%, and $\text{AP}_{50}^{Seal}$ of 98.5%. Therefore, YOLO11m can be considered an effective model for the DKDS dataset.
5.2.2 Visualization
We present the prediction results of various YOLO models for text and seal detection, along with the corresponding input and ground-truth images. As shown in Fig. 10, seals are marked with yellow bounding boxes, while Kuzushiji characters are marked with blue bounding boxes.
In the first row, YOLOv10m failed to detect a Kuzushiji character at the bottom of the third row from the left, while the other models detected all characters correctly. However, all models incorrectly predicted stains in the upper-left corner as Kuzushiji characters, which is the main factor affecting model performance. Furthermore, although our ground-truth annotations do not include positional information for marginal notes, both YOLOv8m and YOLO11m successfully detected several marginal notes, as shown on the far right.
In the second row, the ink on the Kuzushiji character in the upper-left corner is fading, which poses a challenge for detection. None of the YOLO models except YOLO11m correctly detected this character, demonstrating YOLO11m’s superior capability in Kuzushiji character detection.
5.3 Track 2: Document Binarization
5.3.1 Quantitative Comparison
We employed traditional binarization algorithms, traditional algorithms combined with K-means clustering, and deep learning-based (i.e., GAN) methods as baselines for the document binarization task. The experimental results are summarized in Table 3.
Overall, the traditional binarization algorithms presented poor performance. Specifically, Niblack’s method [niblack1985introduction] performed the worst, as indicated by its high DRD value, reflecting strong sensitivity to noise in the DKDS dataset. Otsu’s method [otsu1979threshold], a simple global thresholding method, achieved moderate results, while Sauvola’s method [sauvola2000adaptive] significantly outperformed both by leveraging locally adaptive thresholding to process complex backgrounds more effectively.
Furthermore, applying K-means clustering [mcqueen1967some] as a preprocessing step to remove red interference (e.g., seals) further improved performance. For instance, the ASM increased from 27.25 to 28.49, from 50.10 to 69.57, and from 72.55 to 73.09 for K-means + Niblack, K-means + Otsu, and K-means + Sauvola, respectively.
In contrast, the deep learning–based method (i.e., GAN) demonstrated superior performance, achieving the FM of 98.11, p-FM of 98.14, PSNR of 26.53 dB, DRD of 0.82, and an ASM of 80.49. The high FM, p-FM, and low DRD values indicate that the generated binary images are highly readable. However, the relatively lower PSNR suggests that, although the Kuzushiji character outlines are clear, some noise or imperfections remain along the edges.
5.3.2 Visualization
We present two groups of result images generated by different methods, along with their corresponding input and ground-truth images, as shown in Figs. 11 and 12. The input images contain various forms of degradation, including red seals, paper damage, stains, and ink fading, while the ground-truth images remove all such interference, retaining only the black Kuzushiji characters.
Among the traditional methods, Niblack’s algorithm [niblack1985introduction] produced the poorest results, with much noise remaining in the background. Otsu’s method [otsu1979threshold] achieved better separation between text and background but still retained some noise. Sauvola’s method [sauvola2000adaptive] performed relatively better but failed to completely remove seal interference.
When combined with K-means clustering [mcqueen1967some] preprocessing, all three traditional methods achieved improved performance, further reducing seal marks, though minor residues remained.
In contrast, the GAN-based method [goodfellow2020generative] obtained the best results, effectively removing most noise and seal interference while maintaining excellent text legibility.
6 Conclusion
Binarizing degraded Kuzushiji document images presents significant challenges, particularly when seals overlap with the characters. Effectively removing seal interference while preserving or restoring character details remains a critical problem. Notably, no publicly available dataset currently contains Kuzushiji documents with seals, limiting the development of methods for this task. To address this gap, we constructed the Degraded Kuzushiji Documents with Seals (DKDS) dataset by collecting and integrating pre-modern Kuzushiji document images with high-resolution images of ancient seals. We detail the construction process of the DKDS dataset and define two task tracks based on it: (1) text and seal detection, and (2) document binarization. Furthermore, we provide baseline evaluation results for several classical methods to facilitate future research.
Notably, the selected Kuzushiji document dataset includes corresponding Unicode character mappings and OCR annotations, providing valuable resources for training models on both document binarization and OCR tasks. In future work, we plan to jointly design and train models for these tasks, aiming to develop an end-to-end system capable of efficiently and accurately converting Kuzushiji characters into modern Japanese in degraded pre-modern Japanese documents.
Acknowledgments
We sincerely appreciate Jie Chen for kindly providing and supporting the collection of the raw seal data used in this work.
Funding
This work was supported by JSPS KAKENHI Grant Numbers 25H01242.
Author Contributions
Rui-Yang Ju: Conceptualization, Data Curation, Formal Analysis, Methodology, Writing – Original Draft Preparation, Writing – Review & Editing; Kohei Yamashita: Data Curation, Investigation, Writing – Review & Editing; Hirotaka Kameko: Project Administration, Resources, Writing – Review & Editing; Shinsuke Mori: Funding Acquisition, Resources, Supervision, Writing – Review & Editing.
Data availability
The proposed dataset and implementation code are publicly available on GitHub at https://github.com/RuiyangJu/DKDS.
Declarations
Competing Interests
The authors declare that they have no conflict of interest.
Ethics approval
This research does not involve human participants and/or animals.

### v2

@@
 We describe the dataset construction process, which required the assistance of a trained Kuzushiji expert, and define two benchmark tracks: (1) text and seal detection and (2) document binarization.
-For the text and seal detection track, we provide baseline results using multiple versions of the You Only Look Once (YOLO) models for detecting Kuzushiji characters and seals.
-For the document binarization track, we present baseline results from traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods.
+For the text and seal detection track, we provide baseline results using several recent versions of the You Only Look Once (YOLO) models for detecting Kuzushiji characters and seals.
+For the document binarization track, we present baseline results from traditional binarization algorithms, traditional algorithms combined with K-means clustering, two state-of-the-art (SOTA) Generative Adversarial Network (GAN) methods, as well as our Conditional GAN (cGAN) baseline.
 The DKDS dataset and the implementation code for baseline methods are available at https://ruiyangju.github.io/DKDS.
 keywords:
-Kuzushiji, Seals, Pre-modern Japanese Documents, Benchmark Dataset, Object Detection, Document Binarization
+Kuzushiji, Pre-modern Japanese Documents, Benchmark Dataset, Object Detection, Document Binarization, Seals
+Figure 1: Comparison of Optical Character Recognition (OCR) results on Kuzushiji characters overlapping with seals between the original and binarized images.
+OCR was conducted using the “miwo” app [clanuwat2021miwo].
+From left to right, the observed OCR errors include recognition of extra character, recognition of incorrect character, and misclassification of seal inscriptions as text.
 1 Introduction
-Pre-modern Japanese documents were mainly written in Kuzushiji111くずし字, a form of Japanese cursive script (草書体)., a pre-modern Japanese cursive script widely used in official records, personal letters, and literary works.
+Pre-modern Japanese documents were mainly written in Kuzushiji111くずし字, a form of Japanese cursive script (草書体)., which is a cursive script widely used in official records, personal letters, and literary works.
 Kuzushiji is characterized by complex, flowing strokes and highly variable forms, which differ markedly from modern Japanese characters [ueki2021survey].
@@
 Following the adoption of modern kana orthography in 1946, most contemporary native Japanese speakers are unable to read Kuzushiji texts without specialized training [hashimoto2017Kuzushiji].
-Figure 1: Comparison of Optical Character Recognition (OCR) results on Kuzushiji characters overlapping with seals between the original and binarized images.
-OCR was conducted using the “miwo” app [clanuwat2021miwo].
-From left to right, the observed OCR errors include recognition of extra character, recognition of incorrect character, and misclassification of seal inscriptions as text.
-With the remarkable advances in deep learning for computer vision and natural language processing, researchers have increasingly applied neural network models to Kuzushiji character recognition [le2019human, le2020automated].
-While trained models typically achieve satisfactory performance in recognizing Kuzushiji characters seen during training, their performance often drops significantly when tested on unseen (zero-shot) Kuzushiji characters.
-For instance, the framework proposed by Ishikawa et al.
-[ishikawa2024japanese] achieved an accuracy of only approximately 48% in such zero-shot scenarios.
-This result presents the critical need to construct high-quality Kuzushiji datasets, including special cases, to improve model generalization and capabilities in complex recognition tasks.
-The primary Kuzushiji datasets include Kuzushiji-MNIST, Kuzushiji-49, and Kuzushiji-Kanji [^2] Kanji refers to Chinese-origin characters used in Japanese.
+With the remarkable advances in deep learning for computer vision, researchers have increasingly applied neural network models to Kuzushiji character recognition.
+The primary datasets for Kuzushiji character recognition include Kuzushiji-MNIST, Kuzushiji-49, and Kuzushiji-Kanji [^2] Kanji refers to Chinese-origin characters used in Japanese.
 [clanuwat2018deep].
-The Kaggle competitions organized using these datasets has further advanced both research and practical applications [kitamoto2019progress, kitamoto2020kaggle].
+The Kaggle competitions organized using these datasets have further advanced both research and practical applications [kitamoto2019progress, kitamoto2020kaggle].
 Together, these datasets and competitions cover a wide range of Kuzushiji characters and provide a standardized evaluation platform for researchers.
-With these resources, researchers can perform not only Kuzushiji character classification but also related tasks such as Kuzushiji detection and generation [soan2023], thereby advancing the development of Kuzushiji character recognition techniques and enhancing model performance.
-Several methods have been proposed for the Kuzushiji character recognition task.
-Fujita et al.
-[fujita2024learning] employed Transformer [vaswani2017attention] models to analyze and recognize Kuzushiji characters, leveraging attention mechanisms to capture relationships between adjacent characters and accurately separate character regions and boundaries.
-Ngo et al.
-[ngo2021recurrent] introduced RNN-Transducer [graves2012sequence] models for handwritten Kuzushiji character recognition, effectively exploiting both visual and text information from the input images and achieving satisfactory performance.
+With these resources, researchers can perform not only Kuzushiji character classification but also related tasks such as Kuzushiji detection and generation [soan2023], thereby advancing the development of Kuzushiji character recognition and enhancing model performance.
 Figure 2: DKDS dataset is the first collection of degraded pre-modern Japanese document images specifically designed to address the challenge of Kuzushiji characters overlapping with seals.
 Based on the dataset, we define two benchmark tracks: (1) Text and Seal Detection, and (2) Document Binarization.
-In addition, KuroNet [clanuwat2019kuronet, lamb2020kuronet] is an end-to-end model based on a residual U-Net [ronneberger2015u] architecture, designed to capture long-range context, process extensive vocabularies, and accommodate non-standard character layouts, thereby enabling the recognition of pre-modern Japanese documents.
-On the Kuzushiji dataset provided by the Kaggle competitions [kitamoto2019progress, kitamoto2020kaggle], KuroNet achieved an accuracy exceeding 90%.
-To facilitate the deployment of OCR models on mobile devices, Clanuwat et al.
-[clanuwat2021miwo] developed the “miwo” application, which allows users to upload or photograph Kuzushiji documents and automatically transcribe the characters into modern Japanese.
+Meanwhile, the Center for Open Data in the Humanities (CODH) has developed several Kuzushiji OCR systems that address both character-level and page-level recognition scenarios.
+In 2021, CODH introduced the mobile application “miwo” [clanuwat2021miwo], which enables users to capture or upload images of Kuzushiji documents and automatically transcribe them into modern Japanese.
+CODH also released KuroNet [clanuwat2019kuronet, lamb2020kuronet], an end-to-end OCR model based on a Residual U-Net architecture [ronneberger2015u].
+Designed to capture long-range context, process extensive vocabularies, and accommodate non-standard character layouts, KuroNet supports robust recognition of pre-modern Japanese texts.
+On Kaggle’s Kuzushiji datasets [kitamoto2019progress, kitamoto2020kaggle], KuroNet achieved classification accuracies exceeding 90%.
+More recently, SakanaAI introduced Metom [Metom], a Vision Transformer [dosovitskiy2021vit]-based Kuzushiji classifier specialized for single-character recognition.
+Metom was trained on 2,703 types of Kuzushiji characters that appear at least five times in the large-scale Kuzushiji document provided by CODH [genjimonogatari], offering a powerful and extensible foundation for character-level analysis.
 The aforementioned OCR methods and applications perform well on clean pre-modern Japanese documents written in Kuzushiji; however, they do not sufficiently consider the effects of noise, such as document degradation and seals, on recognition accuracy.
@@
 To address this issue, document binarization, which preserves the foreground (black) while removing background and noise (white), has been widely adopted as a preprocessing step, as it can effectively improve OCR performance.
-Among the various types of noise, the overlap between Kuzushiji characters and seals poses a major challenge.
-Serving as symbols of ownership and collector identity, seals were widely used in ancient and pre-modern Asian document collections [li2021character].
-They are typically inscribed with stylized ancient scripts and may contain not only the collector’s name but also phrases expressing personal wishes, interests, or social status.
-In ancient and pre-modern documents, seals usually appear as red marks that frequently overlap with characters, which can blur the characters and lead to OCR errors.
+Seals, serving as symbols of ownership and collector identity, are widely used in ancient and pre-modern Asian documents [li2021character].
+In these documents, seals usually appear as red marks inscribed with stylized ancient scripts, and may contain the collector’s name as well as phrases reflecting personal wishes, interests, or social status.
+Among the various types of noise, the overlap between Kuzushiji characters and seals poses a major challenge for text and seal detection as well as document binarization.
 To address this challenge, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset, developed in collaboration with a trained Kuzushiji expert.
@@
 2, we define two task tracks for the DKDS dataset: (1) text and seal detection, and (2) document binarization.
-The former track aims to accurately detect text (i.e., Kuzushiji characters) and seal regions within the documents, while the latter focuses on preserving the primary textual content while removing noise, including stains and seals, thereby providing clean input for subsequent OCR tasks.
+The former track aims to accurately detect text (i.e., Kuzushiji characters) and seal regions within the documents, while the latter focuses on preserving the primary textual content while removing noise, including stains and seals, thereby providing clean input for the subsequent OCR task.
 Overall, the contributions of this work are summarized as follows:
@@
     We introduce a novel benchmark dataset, named Degraded Kuzushiji Documents with Seals (DKDS), developed in collaboration with a trained Kuzushiji expert.
-To the best of our knowledge, this is the first publicly available dataset specifically addressing the challenge of Kuzushiji characters overlapped with seals.
+To the best of our knowledge, this is the first publicly available dataset specifically addressing the challenge of Kuzushiji characters overlapping with seals.
   • (b)
@@
     We define the text and seal detection task, which serves as a preliminary step for recognizing Kuzushiji characters and transcribing them into modern Japanese.
-Baseline results for this task are also presented using multiple versions of the YOLO models.
+Baseline results for this task are also presented using several recent versions of the YOLO models.
   • (d)
@@
 Compared to existing binarization datasets, the DKDS dataset presents additional challenges due to the overlap of Kuzushiji characters with seals.
-Baseline results for this task are provided, including traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods.
+Baseline results for this task are provided, including traditional binarization algorithms, traditional algorithms combined with K-means clustering, and two state-of-the-art (SOTA) Generative Adversarial Network (GAN) methods, as well as our Conditional GAN (cGAN) baseline.
 The rest of this paper is organized as follows: Section 2 reviews related datasets and previous work relevant to the two task tracks defined here, highlighting their limitations.
@@
 2.1 Related Dataset
-The datasets relevant to this work can be categorized into three domains: (1) Kuzushiji document datasets, (2) degraded pre-modern document binarization datasets, and (3) seal datasets.
-Existing research on pre-modern Japanese documents written in Kuzushiji [clanuwat2018deep, le2019human] lacks a dataset that specifically addresses the challenge of overlaps between characters and seals, which can significantly reduce recognition accuracy.
-Most publicly available datasets for degraded ancient document binarization focus on documents written in Latin-based alphabetic systems.
-For instance, the documents in the Document Image Binarization Contest (DIBCO) series datasets [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2012icfhr, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019] are primarily written in English, while the Bickley Diary (BD) dataset [deng2010binarizationshop], Synchromedia Multispectral Ancient Document Images (SMADI) [hedjam2013historical], and Persian Heritage Image Binarization Dataset (PHIBD) [ayatollahi2013persian] all contain documents written in European languages.
+The datasets relevant to this work can be categorized into three domains: (1) Kuzushiji document datasets, (2) degraded document binarization datasets, and (3) seal datasets.
+Existing research [clanuwat2018deep] on pre-modern Japanese documents written in Kuzushiji lacks a dataset that specifically addresses the challenge of overlaps between characters and seals, which can significantly reduce recognition accuracy.
+Most publicly available datasets for degraded document binarization focus on documents written in Latin-based alphabetic systems.
+For instance, the documents in the Document Image Binarization Contest (DIBCO) series datasets [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2012icfhr, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019] are primarily written in English, while the Bickley Diary (BD) [deng2010binarizationshop], Synchromedia Multispectral Ancient Document Images (SMADI) [hedjam2013historical], and Persian Heritage Image Binarization Dataset (PHIBD) [ayatollahi2013persian] all contain documents written in alphabetic scripts.
 In contrast, no publicly available document binarization datasets exist for pre-modern Japanese documents, particularly those written in Kuzushiji.
 Furthermore, no existing degraded document datasets include seals as the primary source of interference.
-Therefore, we aim to randomly add seals to the collected pre-modern Japanese documents written in Kuzushiji.
-However, existing publicly available seal datasets are all unsuitable, as the seals they contain are either not ancient Asian seals, suffer from poor image quality, or have complex backgrounds.
-Specifically, the Chinese-Seal Dataset (CSD) [xu2025chinese] focuses on modern Chinese seals; DDI-100 [zharikov2020ddi] contains Western (English-language) seal samples; and SPODS [nandedkar2016spods] comprises seal images from contemporary office settings.
-Although MiikeMineStamps [buitrago2021miikeminestamps] includes some Japanese seal images, they are generally low-resolution and feature non-uniform backgrounds, making background removal difficult.
+Yang et al.
+[yang2023docdiff] synthesized paired data by overlaying seal images onto clean documents; however, their synthesized documents that primarily contain modern Chinese characters, which differ significantly from the pre-modern Japanese Kuzushiji characters targeted in this work.
+Following a similar strategy, we also simulate seal-induced degradation by randomly adding seal images to collected pre-modern Japanese documents.
+Regarding the source of seal images, existing publicly available seal datasets are unsuitable for this purpose, as they either do not contain ancient Asian seals, suffer from poor image quality, or exhibit complex backgrounds.
+For instance, MiikeMineStamps [buitrago2021miikeminestamps] includes some Japanese seal images; however, they are generally low-resolution and feature non-uniform backgrounds, making background removal difficult.
 2.2 Text and Seal Detection
-The diverse shapes of ancient seals, their curved inscriptions, and the frequent occurrence of text overlap pose significant challenges for seal processing and text analysis and recognition.
-As a critical preliminary step for Kuzushiji character recognition, text detection is particularly important [li2023yolo].
+The diverse shapes of ancient seals, their curved inscriptions, and the frequent occurrence of text overlap pose significant challenges for seal processing as well as text analysis and recognition.
+As a critical preliminary step in Kuzushiji character recognition, text detection is particularly important.
+For instance, DBNet [liao2020real], a widely used text detection method in OCR systems, achieved an F-measure of 85.4% at 26 frames per second on the ICDAR 2015 [karatzas2015icdar] dataset, demonstrating real-time performance with high detection accuracy.
 Similarly, seal detection is important for subsequent seal processing.
 However, the scarcity of ancient seal samples and the prevalence of low-quality or degraded images severely limit available training data and annotations for related tasks [li2021character].
-For instance, Micenkov et al.
-[micenkov2011stamp] proposed an automatic seal segmentation system for insurance documents that achieved 84% accuracy.
+Micenkov et al.
+[micenkov2011stamp] proposed an automatic seal segmentation system for insurance documents.
+When evaluated on the collected dataset of 400 document images, named StaVer, their system achieved a recall of 83% and a precision of 84%.
 Yu et al.
 [yu2023icdar] developed a modern seal dataset for official and financial contexts, introducing two tasks: seal title text detection and end-to-end seal title recognition.
-To the best of our knowledge, no dataset currently exists for text and seal detection in degraded pre-modern Japanese Kuzushiji documents.
-The YOLO series of models [jocher2023yolov8, wang2024yolov9, wang2024yolov10, jocher2024yolo11] achieve an excellent balance between detection accuracy and model size, demonstrating outstanding performance across a wide range of object detection tasks [ju2023fracture, chien2025yolov8, chien2024yolov9, ju2024yolov8, ju2024global].
-Researchers have explored the application of the YOLO series of models in Kuzushiji character detection [li2023yolo] and seal detection [bento2025performance].
-For instance, Bento et al.
-[bento2025performance] merged and adapted the StaVer [micenkov2011stamp] and DDI-100 [zharikov2020ddi] datasets to evaluate the performance of YOLOv8, YOLOv9, YOLOv10, and YOLO11 on their newly constructed dataset for seal detection.
-Similarly, Sun et al.
-[sun2024ra] enhanced the YOLOv8 architecture and proposed RA-YOLOv8, achieving higher detection accuracy.
-Nevertheless, due to the lack of seal data based on Kuzushiji documents, these improved YOLO models cannot be directly applied to this specific scenario.
+To the best of our knowledge, no dataset currently exists for text and seal detection in degraded pre-modern Japanese documents written in Kuzushiji.
+The YOLO series of models [jocher2023yolov8, wang2024yolov9, wang2024yolov10, jocher2024yolo11] achieve an excellent balance between detection accuracy and model size, demonstrating outstanding performance across a wide range of object detection tasks.
+Researchers have explored the application of the YOLO series of models in text detection and seal detection.
+Nevertheless, due to the lack of ancient seal data from pre-modern Japanese documents, the recent versions of YOLO models cannot be directly applied to this specific scenario.
+2.3 Document Binarization
+Document binarization is an important preprocessing step for OCR, as it effectively removes various types of noise, such as paper yellowing, text fading, ink bleeding, stains, and seals.
+Early work in this area primarily relied on threshold-based methods.
+Specifically, Otsu [otsu1979threshold] introduced the global thresholding algorithm in 1979.
+Subsequently, Niblack [niblack1985introduction] and Sauvola [sauvola2000adaptive] proposed local adaptive thresholding techniques in 1985 and 2000, respectively, enabling more robust binarization of documents with non-uniform illumination and complex background variations.
+To facilitate systematic evaluation, the Document Image Binarization Contest (DIBCO), one of the most popular competitions in this field, released benchmark datasets from 2009 to 2019 [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2012icfhr, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019].
+These datasets have played an important role in advancing both traditional and deep learning-based binarization methods by providing standardized document image evaluation benchmarks.
+The introduction of GANs [goodfellow2020generative] has enabled the generation of high-quality binary document images.
+Suh et al.
+[suh2022two] proposed a two-stage GAN framework that employed six enhanced CycleGANs [zhu2017unpaired] to produce binarized document images.
+Building on this line of work, Ju et al.
+[ju2023ccdwt, ju2024three] further introduced a three-stage GAN architecture, was also composed of six improved CycleGAN modules, to advance document binarization performance.
+Both methods consistently achieved SOTA results on the DIBCO benchmarks.
+However, neither traditional binarization methods nor SOTA GAN-based methods have been evaluated on degraded pre-modern Japanese documents written in Kuzushiji, particularly those containing seals, overlapping text, and complex degradations.
+This gap highlights the need to develop datasets and models that are tailored to the unique characteristics of pre-modern Japanese documents.
 Figure 3: Examples of raw Kuzushiji document data from the book Genji Monogatari (The Tale of Genji[genjimonogatari].
-2.3 Document Binarization
-Document binarization is an important preprocessing step for OCR and layout analysis, effectively removing various types of noise from document images, such as paper yellowing, text fading, ink bleeding, stains, and seals.
-The Document Image Binarization Contest (DIBCO), one of the most popular competitions in this field, has provided benchmark datasets from 2009 to 2019 [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2012icfhr, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019].
-Researchers have proposed a range of neural network models evaluated on these datasets, including those based on GANs [suh2022two, ju2023ccdwt, ju2024mfe], Transformers [souibgui2022docentr], and diffusion models [saifullah2023coldbin, yang2023docdiff], which have significantly advanced the field.
-However, most existing datasets contain ancient and pre-modern document images based on Latin alphabets, while few include Kuzushiji documents, particularly those containing overlapping seals and text.
-Figure 4: Examples of our collected imperial seals from the Qing dynasty, with backgrounds removed.
-Figure 5: The overall workflow of the proposed DKDS dataset construction includes raw data collection, detection annotations, initial binarization ground-truth generation, verification, and manual correction.
-The initial binarization ground-truth was generated using the model proposed by Ju et al.
-[ju2024three], while the verification was conducted by a trained Kuzushiji expert.
 3 Dataset Construction
 3.1 Raw Data Collection
-The raw Kuzushiji document data were obtained from the Center for Open Data in the Humanities (CODH) [genjimonogatari].
+The raw Kuzushiji document data were obtained from the CODH [genjimonogatari].
 Specifically, we selected the dataset corresponding to book ID “200003803”, titled Genji Monogatari (The Tale of Genji) [^3] https://codh.rois.ac.jp/char-shape/book/200003803/.
-Examples from this book are shown in Fig 3.
+Examples from this book are shown in Fig.
+3.
 This dataset was chosen for two main reasons: (1) its historical significance and widespread popularity; and (2) the availability of detailed OCR annotations provided by CODH, including 237 unique character types and a total of 11,132 characters.
-These data can be used for training and evaluating OCR models in future research.
+These data can also be used for training and evaluating OCR models in future research.
 Furthermore, to date, there is no publicly available high-resolution dataset of ancient Japanese seals.
-To address this gap, we collected 219 seals from three emperors of the Qing dynasty (i.e., Kangxi (1661–1722), Yongzheng (1722–1735), and Qianlong (1735–1796)) sourced from the developer of an ancient seal app [^4] https://www.daizigege.com/seal/.
-The primary reason for selecting ancient Chinese seals is that both ancient Japanese and Chinese seals are mainly inscribed with kanji characters.
+To address this gap, we collected 219 seals belonging to three Chinese ancient emperors of the Qing dynasty (i.e., Kangxi (1661–1722), Yongzheng (1722–1735), and Qianlong (1735–1796)) sourced from the developer of an ancient seal app [^4] https://www.daizigege.com/seal/.
+The primary reason for selecting ancient Chinese seals is that both ancient Japanese and Chinese seals are mainly inscribed with Kanji characters, and no high-resolution images of ancient Japanese seals are currently publicly available.
 Considering factors such as image resolution and seal patterns (e.g., whether the inscriptions are written in Kanji), we selected a final set of 128 seal images from the collection.
@@
 4, these seals are high-resolution, and their backgrounds can be easily removed.
-Table 1: Statistics of the proposed DKDS dataset.
-Category Total Training Testing
-Kuzushiji Documents 50 40 10
-Seals 128 $100\times 2$ 28
-Seals per Document - 4 – 6 2 – 3
-Figure 6: Examples of verification and manual correction in the dataset construction workflow: “Raw” refers to the input images, “Initial” denotes the initial binarization produced by Ju et al.
-[ju2024three], and “Corrected” represents the final ground-truth after verification and manual correction.
-The trained Kuzushiji expert performed verification to determine which elements should be preserved (e.g., marginal notes, voiced sound symbols, periods) and which should be removed (e.g., stains).
-Figure 7: Challenging examples for the text and seal detection task: The first row presents seals suffering from ink fading, while the second row shows seals overlapping with Kuzushiji characters or other seals.
+Figure 4: Examples of our collected imperial seals from the Qing dynasty, with backgrounds removed, which are used to simulate seal interference.
 3.2 Construction Workflow
@@
 5.
-To create the final document images, we combined 128 background removed seal images with 50 Kuzushiji document images.
+To create the final document images, we combined 128 background-removed seal images with 50 Kuzushiji document images.
 Specifically, 40 Kuzushiji documents were selected for training, and the remaining 10 for testing.
 As detailed in Table 1, the 128 seal images were divided into 100 for the 40 training Kuzushiji documents, and 28 for the 10 test Kuzushiji documents.
-To increase the cases of overlap between seals and Kuzushiji characters in the training set, we duplicated the 100 training seal images to create 200 seals, which were then randomly added onto the 40 training Kuzushiji documents.
-As a result, each Kuzushiji document image in the training set contains approximately 4–6 seals, while those in the test set contain about 2–3 seals.
+To increase the cases of overlap between seals and Kuzushiji characters in the training set, we duplicated the 100 training seal images to create 200 seals, which were then randomly added to the 40 training Kuzushiji documents.
+As a result, each Kuzushiji document image in the training set contains approximately 4–6 seals.
+In addition, we constructed easy-level (Testing-E) and difficult-level (Testing-D) versions of the test set according to the number of seals per document.
+Each document image in Testing-E contains approximately 2–3 seals, while each document image in Testing-D contains 14 seals.
 To generate the binarization ground-truth, we employed a combined approach of neural network generation followed by manual verification and correction.
 As shown in Fig.
-5, we first applied the model proposed by Ju et al.
-[ju2024three] to binarize the raw Kuzushiji document images (without added seals), producing the initial binarization ground-truth.
+5, we first applied a pre-trained binarization model trained on the DIBCO benchmarks to binarize the raw Kuzushiji document images (without added seals), producing the initial binarization ground-truth.
 Subsequently, a trained Kuzushiji expert manually verified these initial results.
 Even in the absence of seal interference, the initial binarization often contained errors.
-As detailed in Fig.
+As shown in Fig.
 6, the expert identified issues such as missing marginal notes, voiced sound symbols, periods, and unremoved stains.
@@
 Notably, this verification step was critical and required a trained Kuzushiji expert to ensure the validity of the process.
+Figure 5: The overall workflow of the proposed DKDS dataset construction includes raw data collection, text and seal detection annotations, initial binarization ground-truth generation, verification, and manual correction.
+The initial binarization ground-truth was generated using a pre-trained binarization model trained on the DIBCO benchmarks, while the verification was conducted by a trained Kuzushiji expert.
+Figure 6: Examples of verification and manual correction in the dataset construction: “Raw” refers to the input images, “Initial” denotes the initial binarized images produced by a pre-trained binarization model, and “Corrected” represents the final ground-truth images after verification and manual correction.
+The trained Kuzushiji expert performed verification to determine which elements should be preserved (e.g., marginal notes, voiced sound symbols, and periods) and which should be removed (e.g., stains).
+Table 1: Statistics of the proposed DKDS dataset.
+Category Total Training Testing-E Testing-D
+Document 50 40 10 10
+Seal 128 $100\times 2$ 28 $28\times 5$
+S/D - 4 – 6 2 – 3 14
+S/D denotes the number of seals per document.
 3.3 Detection Annotations
@@
 As shown in Fig.
-7, seals may overlap with Kuzushiji characters or other seals, and suffer from ink fading, all of which can significantly reduce detection accuracy, making this task highly challenging.
+7, seals may overlap with Kuzushiji characters or with other seals, and suffer from ink fading.
+These degradations can lead to incomplete seal detection or significantly reduce Kuzushiji character detection accuracy, making the detection task highly challenging.
 In the document binarization task, the goal is to remove seals while preserving, or even restoring, the underlying Kuzushiji characters.
@@
 Text and seal detection serves as a critical preliminary step for downstream applications, including Kuzushiji character recognition and seal processing, while document binarization aims to improve the accuracy of subsequent OCR systems.
-By addressing these challenges, the proposed dataset encourages the development of robust models capable of processing real-world scenarios in Kuzushiji document analysis.
-Figure 8: Challenging examples for the document binarization task: The left column shows the input Kuzushiji document images, while the right column presents the corresponding binarized images.
+By addressing these challenges, the proposed DKDS dataset encourages the development of robust models capable of processing real-world scenarios in Kuzushiji document analysis.
 4.2 Evaluation Metrics
-For the text and seal detection task, we adopted standard object detection evaluation metrics [lin2014microsoft], including model parameters (Params), floating-point operations (FLOPs), average precision at a 50% intersection over union (IoU) threshold ($\text{AP}_{50}$), and average precision across IoU thresholds from 50% to 95% ($\text{AP}_{50:95}$).
-Model parameters reflect the model’s size and complexity, FLOPs quantify computational cost, and these average precision (AP) metrics provide a comprehensive measure of detection accuracy.
-For the document binarization task, we employed classical evaluation metrics [gatos2009icdar] for quantitative comparison, including the F-measure (FM), pseudo F-measure (p-FM), peak signal-to-noise ratio (PSNR), and distance reciprocal distortion (DRD).
-To provide a more comprehensive assessment, we also adopted the Average-Score Metric (ASM) proposed by Ju et al.
-[ju2025efficient], which is calculated as follows:
-$ASM=\dfrac{FM+p\text{-}FM+PSNR+(100-DRD)}{4}.$ (1)
-A higher ASM value indicates better overall performance, providing a single unified metric for model ranking.
-Figure 9: The architecture of the GAN-based method used.
+For the text and seal detection task (Track 1), we adopt standard evaluation metrics commonly used in text detection [karatzas2015icdar, liao2020real], including the number of model parameters (Params), floating-point operations (FLOPs), precision (P, in %), recall (R, in %), and F-measure (F, in %).
+For the document binarization task (Track 2), we employ classical evaluation metrics [gatos2009icdar] for quantitative comparison, including the F-measure (FM), pseudo F-measure (p-FM), peak signal-to-noise ratio (PSNR), and distance reciprocal distortion (DRD).
+To provide a more comprehensive assessment, we also adopt the Average-Score Metric (ASM) proposed by Ju et al.
+[ju2025efficient, ju2025mfegan], which is calculated as follows:
+$\text{ASM}\!=\!\frac{1}{4}(\text{FM}\!+\!\text{p-FM}\!+\!\text{PSNR}\!+\!(100\!-\!\text{DRD)}),$ (1)
+where a higher ASM value indicates better overall performance, providing a single unified metric.
 4.3 Baselines
-For the text and seal detection task, we employed the YOLO series of object detection models, including YOLOv8 [jocher2023yolov8], YOLOv9 [wang2024yolov9], YOLOv10 [wang2024yolov10], and YOLO11 [jocher2024yolo11].
+For the text and seal detection task (Track 1), we employ the YOLO series of object detection models, including YOLOv8 [jocher2023yolov8], YOLOv9 [wang2024yolov9], YOLOv10 [wang2024yolov10], and YOLO11 [jocher2024yolo11].
 As one-stage detectors, these models reduce computational complexity and model parameters while maintaining high detection accuracy, achieving a good balance between inference speed and performance.
 In addition, they support end-to-end training, provide fast inference, and present strong generalization capabilities, making them widely adopted in object detection tasks.
-For the document binarization task, we first applied several traditional binarization algorithms to the Kuzushiji documents, including Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive], which can be directly employed without any training.
-Considering that seals in the pre-modern Japanese documents are typically red, while Kuzushiji characters are black and the paper background is pale yellow, we introduced a preprocessing step to remove seal regions based on K-Means clustering [mcqueen1967some] before applying the binarization algorithms.
+For the document binarization task (Track 2), we apply several traditional binarization algorithms to the Kuzushiji documents, including Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive], which can be directly employed without any training.
+Considering that seals in the pre-modern Japanese documents are typically red, while Kuzushiji characters are black and the paper background is pale yellow, we introduce a preprocessing step to remove seal regions based on K-Means clustering [mcqueen1967some] before applying the binarization algorithms.
 This preprocessing effectively reduces the interference caused by red seals and enhances the quality of subsequent binarization results.
-Specifically, we designed and implemented three combined approaches: K-Means + Otsu, K-Means + Niblack, and K-Means + Sauvola.
-In addition, to provide a deep learning-based baseline, we employed a GAN [goodfellow2020generative]-based model to generate binarized Kuzushiji document images.
+Specifically, we implement three combined approaches: K-Means + Otsu, K-Means + Niblack, and K-Means + Sauvola.
+To provide deep learning-based baselines, we employ two SOTA GAN-based methods (i.e., Suh et al. [suh2022two] and Ju et al. [ju2024three]) to evaluate the proposed DKDS dataset.
+In addition, we employ a cGAN [isola2017image]-based method to generate binarized Kuzushiji document images.
 The network architecture for this model is shown in Fig.
 9.
-Since directly inputting the entire image (approximately $2100\times 3200$ pixels) would require excessive GPU memory, we divided each document image into multiple $512\times 512$ patches; the details of this division are described in Section 5.1.2.
-These patches were then fed into a U-Net [ronneberger2015u] architecture with an EfficientNet-B5 [tan2019rethinking] backbone, which served as the generator.
-A simplified PatchGAN [isola2017image, zhu2017unpaired] discriminator was adopted to distinguish between real and fake (generated) images.
+Since directly inputting the entire image (approximately $2100\times 3200$ pixels) would require excessive GPU memory, we divide each document image into multiple $512\times 512$ patches; the details of this division are described in Section 5.1.2.
+These patches are then fed into a U-Net [ronneberger2015u] architecture with an EfficientNet-B5 [tan2019rethinking] backbone, which serves as the generator.
+A simplified PatchGAN [isola2017image, zhu2017unpaired] discriminator is adopted to distinguish between real and fake (generated) images.
 Furthermore, following Suh et al.
 [suh2022two] and Ju et al.
-[ju2024three], we employed a loss function that combines the Wasserstein GAN with Gradient Penalty (WGAN-GP) loss [gulrajani2017improved] and an additional Binary Cross-Entropy (BCE) loss.
-Table 2: Quantitative comparison of various YOLO models for text and seal detection on the DKDS dataset.
-Model Params FLOPs $\text{AP}_{50}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Kuzushiji}}$ $\uparrow$ $\text{AP}_{50}^{\text{Seal}}$ $\uparrow$ $\text{AP}_{50:95}^{\text{Seal}}$ $\uparrow$
-YOLOv8m [jocher2023yolov8] 25.86M 79.1G 96.4% 71.2% 99.1% 86.2%
-YOLOv9m [wang2024yolov9] 20.16M 77.5G 96.3% 71.7% 97.2% 81.4%
-YOLOv10m [wang2024yolov10] 16.49M 64.0G 96.2% 71.4% 99.1% 85.7%
-YOLO11m [jocher2024yolo11] 20.05M 68.2G 97.8% 74.1% 98.5% 85.7%
-Figure 10: Visual comparison of text and seal detection results predicted by different models with the corresponding ground-truth annotations.
-The input images are “200003803_00028_1” (bottom) and “200003803_00028_2” (top).
-Table 3: Quantitative comparison of various methods for document binarization on the DKDS dataset.
-[…]
-Niblack [niblack1985introduction] 39.13 41.14 8.44dB 79.70 27.25
-Otsu [otsu1979threshold] 63.01 63.31 11.76dB 37.69 50.10
-Sauvola [sauvola2000adaptive] 87.87 90.99 18.34dB 7.01 72.55
-K-means [mcqueen1967some] + Niblack [niblack1985introduction] 39.99 42.03 8.61dB 76.67 28.49
-K-means [mcqueen1967some] + Otsu [otsu1979threshold] 84.76 86.28 17.14dB 9.90 69.57
-K-means [mcqueen1967some] + Sauvola [sauvola2000adaptive] 88.59 91.48 18.65dB 6.37 73.09
-GAN [goodfellow2020generative] 98.11 98.14 26.53dB 0.82 80.49
-Figure 11: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00028_2” from the DKDS test set.
-The input image, ground-truth, and method outputs are shown for comparison.
-Figure 12: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00029_1” from the DKDS test set.
-The input image, ground-truth, and method outputs are shown for comparison.
+[ju2024three], we employ a loss function that combines the Wasserstein GAN with Gradient Penalty (WGAN-GP) loss [gulrajani2017improved] and an additional Binary Cross-Entropy (BCE) loss.
+Figure 7: Challenging examples in text and seal detection.
+The first row presents seals suffering from ink fading, while the second row shows seals overlapping with Kuzushiji characters or other seals.
+Figure 8: Challenging examples in document binarization, where Kuzushiji characters overlap with seals.
+Figure 9: The architecture of the cGAN-based method used.
 5 Experiments
 5.1 Implementation Details
-To ensure a fair performance comparison on the DKDS dataset, all models were trained and evaluated on the same hardware (i.e., NVIDIA RTX 3090 GPU).
+To ensure a fair performance comparison on the DKDS dataset, all models were trained and evaluated on the same hardware (i.e., NVIDIA GeForce RTX 3090 GPU).
 All experiments were conducted using the Ubuntu operating system, Python 3.9, and the PyTorch framework.
 5.1.1 Text and Seal Detection
-For the text and seal detection task, the YOLO models were trained for 100 epochs with a batch size of 16 and an input resolution of $640\times 640$ pixels.
+For the text and seal detection task (Track 1), the YOLO models were trained for 100 epochs with a batch size of 16 and an input resolution of $640\times 640$ pixels.
 Model training was performed using the SGD optimizer [robbins1951stochastic] with an initial learning rate of 0.01.
 All YOLO models were initialized with weights pre-trained on the COCO dataset [lin2014microsoft].
+Table 2: Quantitative comparison of medium-sized YOLO models for text and seal detection on the DKDS Testing-E set.
+Model Params FLOPs Kuzushiji Seal
+P R F P R F
+YOLOv8 (Ultralytics) [jocher2023yolov8] 25.86M 79.1G 95.5 89.1 92.2 95.6 96.4 96.0
+YOLOv9 (ECCV2024) [wang2024yolov9] 20.16M 77.5G 95.9 89.9 92.8 99.9 96.4 98.1
+YOLOv10 (NeurIPS2024) [wang2024yolov10] 16.49M 64.0G 93.5 89.9 91.7 99.6 96.4 98.0
+YOLO11 (Ultralytics) [jocher2024yolo11] 20.05M 68.2G 97.1 92.4 94.7 97.0 92.9 94.9
+Table 3: Quantitative comparison of medium-sized YOLO models for text and seal detection on the DKDS Testing-D set.
+Model Params FLOPs Kuzushiji Seal
+P R F P R F
+YOLOv8 (Ultralytics) [jocher2023yolov8] 25.86M 79.1G 90.5 87.9 89.2 92.9 81.4 86.8
+YOLOv9 (ECCV2024) [wang2024yolov9] 20.16M 77.5G 94.0 87.1 90.4 88.4 81.4 84.8
+YOLOv10 (NeurIPS2024) [wang2024yolov10] 16.49M 64.0G 90.4 87.1 88.7 96.4 77.1 85.7
+YOLO11 (Ultralytics) [jocher2024yolo11] 20.05M 68.2G 94.7 89.1 91.8 98.5 84.3 90.8
+Figure 10: Visual comparison of text and seal detection results predicted by different models, together with the corresponding ground-truth annotations.
+The input image is “200003803_00028_2” from the DKDS Testing-D set.
 5.1.2 Document Binarization
-For the document binarization task, traditional binarization algorithms (i.e., Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive]) first convert the input image into a grayscale image as a preprocessing step.
+For the document binarization task (Track 2), traditional binarization algorithms (i.e., Niblack [niblack1985introduction], Otsu [otsu1979threshold], and Sauvola [sauvola2000adaptive]) first convert the input images into grayscale as a preprocessing step.
 The Niblack algorithm uses a $25\times 25$ sliding window with $k=0.8$, while the Sauvola algorithm employs the same window size.
 For the K-Means clustering method [mcqueen1967some], the number of clusters is set to $K=3$.
-After clustering, median filtering with a kernel size of 5 is applied to smooth the results and remove isolated noise points.
-For the GAN-based method [goodfellow2020generative], the training data were preprocessed by dividing each original image into $512\times 512$ overlapping patches with a 30% overlap ratio.
-To further expand the training set, we applied data augmentation, including scaling (factors 0.75, 1.0, 1.25, and 1.5) and rotations (0°and 270°).
-After preprocessing and augmentation, the 40 original training images were expanded into a total of 71,458 patches of size $512\times 512$ pixels.
-Both the generator and discriminator were optimized using the Adam [kingma2014adam] optimizer with a learning rate of $2\times 10^{-4}$ and $\beta$ coefficients of $(0.5,0.999)$.
-The model was trained for 10 epochs with a batch size of 16.
-At inference, the predicted outputs were binarized using a 0.5 threshold.
+After clustering, median filtering with a kernel size of $5\times 5$ is applied to smooth the results and remove isolated noise points.
+For the two SOTA GAN-based methods, we follow their original training data preprocessing strategies, resulting in 325,608 patches of size $256\times 256$ pixels and 240 downscaled full-page images of $512\times 512$ pixels.
+In terms of model architecture, we adopt a U-Net [ronneberger2015u] generator for the method of Suh et al.
+[suh2022two], and employ a U-Net++ [zhou2019unet++] generator for the method of Ju et al.
+[ju2024three].
+Both methods use EfficientNet-B5 [tan2019rethinking] as the encoder backbone.
+For our cGAN [isola2017image]-based method, the training data are preprocessed by dividing each original image into $512\times 512$ overlapping patches with a 30% overlap ratio.
+To further expand the training set, we apply data augmentation, including scaling (factors 0.75, 1.0, 1.25, and 1.5) and rotations (0°and 270°).
+After preprocessing and augmentation, the 40 original training images are expanded into a total of 71,458 patches of size $512\times 512$ pixels.
+Both the generator and discriminator are optimized using the Adam [kingma2014adam] optimizer with a learning rate of $2\times 10^{-4}$ and $\beta$ coefficients of $(0.5,0.999)$.
+The model is trained for 10 epochs with a batch size of 16.
+At inference, the predicted outputs are binarized using a threshold of 0.5.
 5.2 Track 1: Text and Seal Detection
 5.2.1 Quantitative Comparison
-We trained and evaluated various YOLO series models on our proposed DKDS dataset, using the medium (M) size variant for all models. Quantitative comparison results are summarized in Table 2.
-For the text (i.e., Kuzushiji character) detection, all models achieved over 96% in $\text{AP}_{50}$, indicating that they can accurately detect most Kuzushiji characters.
-Under stricter IoU thresholds (i.e., $\text{AP}_{50:95}$), YOLO11m achieved the highest score of 74.1%, outperforming the other models by about 2–3%.
-This demonstrates the superior accuracy of YOLO11m in localizing Kuzushiji character bounding boxes.
-For the seal detection, both YOLOv8m and YOLOv10m obtained the highest $\text{AP}_{50}$ of 99.1%, indicating near-perfect detection performance under the loose IoU criterion.
-In addition, YOLOv8m reached the highest $\text{AP}_{50:95}$ of 86.2%, demonstrating strong localization accuracy under stricter IoU thresholds.
-In terms of model size, YOLOv10m presents the lowest Params of 16.49M and FLOPs of 64.0G, making it the most lightweight model.
-Meanwhile, YOLO11m achieves the best balance between accuracy and efficiency, with 20.05M parameters, 68.2G FLOPs, $\text{AP}_{50}^{Kuzushiji}$ of 97.8%, and $\text{AP}_{50}^{Seal}$ of 98.5%.
-Therefore, YOLO11m can be considered an effective model for the DKDS dataset.
+We train and evaluate various YOLO series models on the proposed DKDS dataset, using the medium (M) size variant for all models, and the quantitative comparison results are shown in Tables 2 and 3.
+On the Testing-E set, all models demonstrate strong overall performance.
+For Kuzushiji text detection, YOLOv11 achieves the best results with an F-measure of 94.7, maintaining an excellent balance between a precision of 97.1 and a recall of 92.4, and significantly outperforms YOLOv8, YOLOv9, and YOLOv10.
+In contrast, for seal detection, YOLOv9 achieves the highest performance with an F-measure of 98.1, closely followed by YOLOv10 with an F-measure of 98.0 and YOLOv8 with an F-measure of 96.0, while YOLOv11 exhibits comparatively lower performance with an F-measure of 94.9.
+The Testing-D set contains a larger number of seals per document, which markedly increases the overall detection difficulty.
+For Kuzushiji text detection, YOLOv11 continues to achieve the highest F-measure of 91.8, demonstrating robust text modeling and detection capability even under complex degradation conditions.
+For seal detection, YOLOv11 shows an even more pronounced advantage, attaining an F-measure of 90.8 and clearly outperforming other models, including YOLOv8 with an F-measure of 86.8, YOLOv9 with 84.8, and YOLOv10 with 85.7.
+Compared with Testing-E, the recall of seal detection decreases markedly for all models on Testing-D, indicating that incomplete seal detection is the primary source of the increased difficulty observed in Testing-D.
+Regarding model efficiency, YOLOv10 has the smallest parameter count at 16.49M and the lowest computational cost at 64.0G FLOPs.
+Meanwhile, YOLOv11 demonstrates an excellent balance between efficiency and performance while maintaining high detection accuracy.
 5.2.2 Visualization
@@
 10, seals are marked with yellow bounding boxes, while Kuzushiji characters are marked with blue bounding boxes.
-In the first row, YOLOv10m failed to detect a Kuzushiji character at the bottom of the third row from the left, while the other models detected all characters correctly.
-However, all models incorrectly predicted stains in the upper-left corner as Kuzushiji characters, which is the main factor affecting model performance.
-Furthermore, although our ground-truth annotations do not include positional information for marginal notes, both YOLOv8m and YOLO11m successfully detected several marginal notes, as shown on the far right.
-In the second row, the ink on the Kuzushiji character in the upper-left corner is fading, which poses a challenge for detection.
-None of the YOLO models except YOLO11m correctly detected this character, demonstrating YOLO11m’s superior capability in Kuzushiji character detection.
+It can be seen that YOLOv10m fails to detect a Kuzushiji character located at the bottom of the third column from the left, whereas the other models successfully detect all characters.
+In addition, all YOLO models incorrectly predicted stains in the upper-left corner as Kuzushiji characters, which constitutes a primary factor affecting detection performance.
+Furthermore, although the ground-truth annotations do not include positional information for marginal notes, all YOLO models detect several marginal notes to varying extents, as shown on the far right of the figure.
+Table 4: Quantitative comparison of various methods for document binarization on the DKDS Testing-E set.
+[…]
+Otsu [otsu1979threshold] 63.01 63.31 11.76dB 37.69 50.10
+Niblack [niblack1985introduction] 39.13 41.14 8.44dB 79.70 27.25
+Sauvola [sauvola2000adaptive] 87.87 90.99 18.34dB 7.01 72.55
+K-means [mcqueen1967some] + Otsu [otsu1979threshold] 84.76 86.28 17.14dB 9.90 69.57
+K-means [mcqueen1967some] + Niblack [niblack1985introduction] 39.99 42.03 8.61dB 76.67 28.49
+K-means [mcqueen1967some] + Sauvola [sauvola2000adaptive] 88.59 91.48 18.65dB 6.37 73.09
+Suh et al. (PR2022) [suh2022two] 187.30M 93.09 93.09 20.99dB 3.12 76.01
+Ju et al. (KBS2024) [ju2024three] 191.46M 95.80 95.81 23.14dB 1.76 78.25
+Ours 31.22M 98.11 98.14 26.53dB 0.82 80.49
+Table 5: Quantitative comparison of various methods for document binarization on the DKDS Testing-D set.
+[…]
+Otsu [otsu1979threshold] 60.51 60.80 11.39dB 40.89 47.95
+Niblack [niblack1985introduction] 37.52 39.41 8.28dB 82.29 25.73
+Sauvola [sauvola2000adaptive] 84.00 87.04 17.08dB 9.60 69.63
+K-means [mcqueen1967some] + Otsu [otsu1979threshold] 68.10 69.04 13.37dB 33.32 54.30
+K-means [mcqueen1967some] + Niblack [niblack1985introduction] 35.98 37.04 8.27dB 82.66 24.66
+K-means [mcqueen1967some] + Sauvola [sauvola2000adaptive] 70.50 69.98 15.80dB 14.69 60.40
+Suh et al. (PR2022) [suh2022two] 187.30M 91.95 91.96 20.10dB 3.85 75.04
+Ju et al. (KBS2024) [ju2024three] 191.46M 94.86 94.90 22.11dB 2.27 77.40
+Ours 31.22M 97.08 97.13 24.58dB 1.38 79.35
+Figure 11: Visual comparison of results generated by different methods on Kuzushiji document “200003803_00027_1” from the DKDS Testing-D set.
+The input image, ground-truth image, and various method outputs are shown for comparison.
 5.3 Track 2: Document Binarization
 5.3.1 Quantitative Comparison
-We employed traditional binarization algorithms, traditional algorithms combined with K-means clustering, and deep learning-based (i.e., GAN) methods as baselines for the document binarization task.
-The experimental results are summarized in Table 3.
-Overall, the traditional binarization algorithms presented poor performance.
-Specifically, Niblack’s method [niblack1985introduction] performed the worst, as indicated by its high DRD value, reflecting strong sensitivity to noise in the DKDS dataset.
-Otsu’s method [otsu1979threshold], a simple global thresholding method, achieved moderate results, while Sauvola’s method [sauvola2000adaptive] significantly outperformed both by leveraging locally adaptive thresholding to process complex backgrounds more effectively.
-Furthermore, applying K-means clustering [mcqueen1967some] as a preprocessing step to remove red interference (e.g., seals) further improved performance.
-For instance, the ASM increased from 27.25 to 28.49, from 50.10 to 69.57, and from 72.55 to 73.09 for K-means + Niblack, K-means + Otsu, and K-means + Sauvola, respectively.
-In contrast, the deep learning–based method (i.e., GAN) demonstrated superior performance, achieving the FM of 98.11, p-FM of 98.14, PSNR of 26.53 dB, DRD of 0.82, and an ASM of 80.49.
-The high FM, p-FM, and low DRD values indicate that the generated binary images are highly readable.
-However, the relatively lower PSNR suggests that, although the Kuzushiji character outlines are clear, some noise or imperfections remain along the edges.
+We employ traditional binarization algorithms, traditional algorithms combined with K-means clustering, and deep learning-based (i.e., GAN) methods as baselines for the document binarization task.
+The experimental results are summarized in Tables 4 and 5.
+On the Testing-E set, the traditional binarization algorithms exhibit generally poor performance.
+Specifically, Niblack’s method [niblack1985introduction] performs the worst, as indicated by its high DRD value, reflecting strong sensitivity to noise in the DKDS dataset.
+Otsu’s method [otsu1979threshold], a simple global thresholding method, achieves moderate results, while Sauvola’s method [sauvola2000adaptive] significantly outperforms both by leveraging locally adaptive thresholding to process complex backgrounds more effectively.
+Furthermore, applying K-means clustering [mcqueen1967some] as a preprocessing step to remove red interference (e.g., seals) further improves performance.
+For instance, the ASM increases from 27.25 to 28.49, from 50.10 to 69.57, and from 72.55 to 73.09 for K-means + Niblack, K-means + Otsu, and K-means + Sauvola, respectively.
+These results indicate that when the number of seals is limited, color-based clustering effectively mitigates seal interference for traditional thresholding methods.
+However, when traditional methods are applied to the Testing-D set, which contains a larger number of seals, the limitations of combining K-means clustering with thresholding methods become more evident.
+Although K-means preprocessing still provides some improvement for simple thresholding approaches, overall performance deteriorates significantly.
+In particular, the FM, p-FM, and DRD metrics for both K-means + Otsu and K-means + Sauvola exhibit noticeable degradation.
+This suggests that in more complex scenarios, clustering-based binarization strategies struggle to robustly process the effects of severe degradation.
+In contrast, deep learning-based methods consistently demonstrate superior performance over traditional methods.
+On the Testing-E set, the method of Suh et al.
+[suh2022two] achieves both FM and p-FM values of 93.09 with a DRD of 3.12, while the method of Ju et al.
+[ju2024three] further improves these results to FM and p-FM values of 95.80 and 95.81, and reduces the DRD to 1.76.
+Our method achieves the best overall performance across all evaluation metrics, attaining FM and p-FM values of 98.11 and 98.14, a PSNR of 26.53 dB, a DRD of 0.82, and the highest ASM of 80.49.
+When evaluated on the more challenging Testing-D set, deep learning-based methods maintain strong robustness.
+The FM score of the method of Suh et al.
+[suh2022two] decreases from 93.09 to 91.95, while the method of Ju et al.
+[ju2024three] exhibits a reduction from 95.80 to 94.86.
+In comparison, our method experiences only a slight decrease from 98.11 to 97.08, while still achieving the lowest DRD value of 1.38.
+Notably, compared with the method of Ju et al.
+[ju2024three], our method further reduces the DRD on Testing-D by approximately 39%, decreasing it from 2.27 to 1.38, while consistently outperforming competing methods in FM, p-FM, and PSNR.
+Finally, in terms of model complexity, the methods of Suh et al.
+[suh2022two] and Ju et al.
+[ju2024three] contain 187.30M and 191.46M parameters, respectively, whereas our model requires only 31.22M parameters, achieving superior performance with a substantially reduced model size.
 5.3.2 Visualization
-We present two groups of result images generated by different methods, along with their corresponding input and ground-truth images, as shown in Figs.
-11 and 12.
+We present visualized results generated by different methods, along with their corresponding input and ground-truth images, as shown in Fig.
+11.
 The input images contain various forms of degradation, including red seals, paper damage, stains, and ink fading, while the ground-truth images remove all such interference, retaining only the black Kuzushiji characters.
-Among the traditional methods, Niblack’s algorithm [niblack1985introduction] produced the poorest results, with much noise remaining in the background.
-Otsu’s method [otsu1979threshold] achieved better separation between text and background but still retained some noise.
-Sauvola’s method [sauvola2000adaptive] performed relatively better but failed to completely remove seal interference.
-When combined with K-means clustering [mcqueen1967some] preprocessing, all three traditional methods achieved improved performance, further reducing seal marks, though minor residues remained.
-In contrast, the GAN-based method [goodfellow2020generative] obtained the best results, effectively removing most noise and seal interference while maintaining excellent text legibility.
+Among the traditional methods, Niblack’s algorithm [niblack1985introduction] produces the poorest results, with much noise remaining in the background.
+Otsu’s method [otsu1979threshold] achieves better separation between text and background but still retains some noise.
+Sauvola’s method [sauvola2000adaptive] performs relatively better but failed to completely remove seal interference.
+Furthermore, as shown in Table 5, the incorporation of K-means clustering does not lead to a substantial performance improvement when documents contain a large number of seals.
+In contrast, deep learning-based methods, including those proposed by Suh et al.
+[suh2022two], Ju et al.
+[ju2024three], and our method, achieve markedly better results by effectively suppressing noise and seal interference while preserving high text legibility.
+5.4 OCR Evaluation
+To quantitatively assess OCR performance on degraded and binarized documents, we employ the “miwo” app to recognize Kuzushiji characters.
+OCR is conducted on combined sets consisting of the training set and each version of the test set, with each evaluation involving a total of 50 document images.
+The experimental results show that binarization leads to a substantial reduction in character error rate (CER).
+On the combined set of 50 images from the training set and the Testing-E set, the CER decreases by 27.6% in relative terms after binarization.
+Similarly, on the combined set of 50 images from the training set and the Testing-D set, the CER is reduced by 34.9%.
+Specifically, for the training set and Testing-E set, the CER of the binarized documents is 1.497%, compared with 2.067% for the original degraded documents.
+For the training set and Testing-D set, the CER of the original degraded documents reaches 2.301%, which is also markedly reduced after binarization (CER = 1.497%).
+Notably, both the raw Kuzushiji document images [genjimonogatari] and the “miwo” OCR app [clanuwat2021miwo] are provided by the Center for Open Data in the Humanities (CODH), which may partially account for the strong recognition performance observed on these documents.
 6 Conclusion
@@
 We detail the construction process of the DKDS dataset and define two task tracks based on it: (1) text and seal detection, and (2) document binarization.
-Furthermore, we provide baseline evaluation results for several classical methods to facilitate future research.
+Furthermore, we provide baseline evaluation results for several classical and deep learning-based methods to facilitate future research.
 Notably, the selected Kuzushiji document dataset includes corresponding Unicode character mappings and OCR annotations, providing valuable resources for training models on both document binarization and OCR tasks.
@@
 Funding
-This work was supported by JSPS KAKENHI Grant Numbers 25H01242.
+This work was supported by JSPS KAKENHI Grant Number 25H01242, and JST SPRING Grant Number JPMJSP2110.
 Author Contributions
-Rui-Yang Ju: Conceptualization, Data Curation, Formal Analysis, Methodology, Writing – Original Draft Preparation, Writing – Review & Editing; Kohei Yamashita: Data Curation, Investigation, Writing – Review & Editing; Hirotaka Kameko: Project Administration, Resources, Writing – Review & Editing; Shinsuke Mori: Funding Acquisition, Resources, Supervision, Writing – Review & Editing.
+Rui-Yang Ju: Conceptualization, Data Curation, Formal Analysis, Methodology, Writing – Original Draft Preparation, Writing – Review & Editing; Kohei Yamashita: Methodology, Data Curation, Investigation, Writing – Review & Editing; Hirotaka Kameko: Project Administration, Resources, Writing – Review & Editing; Shinsuke Mori: Funding Acquisition, Resources, Supervision, Writing – Review & Editing.
 Data availability